# Agent Security Bench (ASB): Formalizing and Benchmarking Attacks and Defenses in LLM-based Agents

### v1

@@ A.3.1 Defenses for Direct Prompt Injection Attack
Delimiters (Learn Prompting, 2023a; Mattern et al., 2023; Willison, 2022).
DPI attacks exploit the agent’s inability to distinguish between user and attacker instructions, leading it to follow the injected prompt. To counter this, we enclose the user instruction within $<$start$>$ and $<$end$>$ delimiters, ensuring the agent prioritizes the user input and ignores the attacker’s instructions.
@@
A.3.2 Defenses for Observation Prompt Injection Attack
Delimiters. Like DPI attacks, OPI attacks exploit the agent’s inability to distinguish between the tool response and the attacker’s instruction, leading it to follow the injected instruction instead of the intended prompt. Therefore, we use the same delimiters as the DPI’s to defend against OPI attacks. The difference is that in OPI the malicious instruction is appended to the observation rather than the user query. Formally, it is defined as:
$\mathbb{E}_{q\sim\pi_{q}}\left[\mathbb{P}(\text{Agent}(\langle\text{start}\rangle q\langle\text{end}\rangle)=a_{b})\right]\gg\mathbb{E}_{q\sim\pi_{q}}\left[\mathbb{P}(\text{Agent}(q\oplus\delta)=a_{m})\right].$ (14)
@@ A.3.2 Defenses for Observation Prompt Injection Attack
Sandwich Prevention (Learn Prompting, 2023c). Since the attack instruction is injected by the tool response during the execution in OPI, the defense method creates an additional prompt and attaches it to the tool response. This can reinforce the agent’s focus on the intended task and redirect its context back, should injected instructions in compromised data have altered it.
@@
A.3.3 Defense for Memory Poisoning Attack
@@
A.3.4 Defenses for PoT Backdoor Attack
@@ C.2 Implementation Details
We conduct all the experiments on the ASB benchmark we designed, which is detailedly introduced in App. B. To ensure the stability of the benchmark results, we conduct all evaluations using simulated tool calls rather than real-world APIs. The DPI, OPI, and Memory Poisoning are experimented with across five prompt injection types in Tab. 1. The mixed and PoT backdoor attacks are tested using combined attacks. PoT attack uses the trigger “with perspicacious discernment” by default.
@@
We implement the LLM-based agent based on the AIOS repository (Mei et al., 2024). For the memory poisoning attack, we focused on using the Prompt Injection Attack to inject into the memory database, as this attack showed the most effective results. The specific LLM used for the injection process was GPT4o-mini. Moreover, the toolset for task execution denoted as $T=T^{t}$, remained consistent throughout the experiments. In addition, we utilized the Chroma vector database [^1] https://python.langchain.com/docs/integrations/vectorstores/ in Langchain to manage and store vectorized representations of agent execution history.
@@
For testing the PoT backdoor attack, we selected five agents from Tab. 8: medical_advisor_agent, legal_consultant_agent, financial_analyst_agent, academic_search_agent, and system_admin_agent. For each agent in the PoT attack, two demonstrative tasks were chosen for ICL. Moreover, two distinct testing tasks were chosen, which are different from the two tasks in the PoT demonstration to ensure the independence and diversity of the experimental results.
C.2.1 Specific Prompts Used
@@
Prompt for Refusal Judgment for Refusal Rate in Sec. 5.2 Following Xie et al. (2024).
@@ D.1 Benchmarking Attacks
Comparing Different Prompt Injection Ways. Tab. 14 displays experimental results for different types of prompt injection types introduced in Tab. 1. In DPI, the Combined Attack excels with a 78.38% ASR by integrating multiple tactics to exploit agent vulnerabilities and obscure malicious intents. In OPI, the Naive attack, with a 28.04% ASR, successfully bypasses defenses due to its simplicity, suggesting that detection mechanisms might be underdeveloped for simpler manipulations. For Memory Poisoning, the Context Ignoring attack leads with an 8.52% ASR by subtly distorting the memory retrieval process without directly altering content.
@@
Attack Type DPI OPI Memory Poisoning Average
@@
Comparisons for Aggressive and Non-aggressive Tasks. Tab. 15 compares the ASR and Refuse Rate for aggressive and non-aggressive tasks. The average ASR for non-aggressive tasks is 38.98%, compared to 33.12% for aggressive tasks, indicating the agent is more vulnerable to attacks on non-aggressive tasks. A possible reason for this is the higher refuse rate for aggressive tasks, which averages 8.31% compared to 4.87% for non-aggressive tasks. This higher refusal rate for aggressive inputs likely helps the agent mitigate more attacks in those scenarios, leading to a lower ASR.

### v2

@@ A.4.1 Defenses for Direct Prompt Injection Attack
Delimiters (Learn Prompting, 2023a; Mattern et al., 2023; Willison, 2022). DPI attacks exploit the agent’s inability to distinguish between user and attacker instructions, leading it to follow the injected prompt. To counter this, we enclose the user instruction within $<$start$>$ and $<$end$>$ delimiters, ensuring the agent prioritizes the user input and ignores the attacker’s instructions.
@@
Dynamic Prompt Rewriting. Since DPI attacks the agent by attaching the attack instruction on the user prompt, Dynamic Prompt Rewriting (DPR) defends against prompt injection attacks by transforming the user’s input query to ensure it aligns with predefined objectives such as security, task relevance, and contextual consistency. The process modifies the original user query $q\oplus\delta$ (where $\delta$ represents the injected malicious instruction) to produce a rewritten query $q^{\prime}=f_{p}(q\oplus\delta)$ using the dynamic rewriting function $f_{p}$.
Unlike paraphrasing, which only alters the sentence structure without changing the underlying meaning, DPR not only changes the sentence structure but also modifies the original meaning to enhance safety. This adjustment reduces the connection between the malicious instruction and the adversarial target action $a_{m}$, making it harder for the injected prompt to influence the agent’s behavior.
The goal is to make the probability of executing the malicious action after dynamic rewriting significantly smaller than without the defense:
Here, $\pi_{q}$ represents the distribution of possible user queries. By dynamically rewriting the prompt, the defense not only removes the potential for malicious content but also alters the meaning in a way that prioritizes safer outcomes for the agent.
A.4.2 Defenses for Indirect Prompt Injection Attack
Delimiters. Like DPI attacks, IPI attacks exploit the agent’s inability to distinguish between the tool response and the attacker’s instruction, leading it to follow the injected instruction instead of the intended prompt. Therefore, we use the same delimiters as the DPI’s to defend against IPI attacks as defined in Eq. 12.
@@ A.4.2 Defenses for Indirect Prompt Injection Attack
Sandwich Prevention (Learn Prompting, 2023c). Since the attack instruction is injected by the tool response during the execution in IPI, the defense method creates an additional prompt and attaches it to the tool response. This can reinforce the agent’s focus on the intended task and redirect its context back, should injected instructions in compromised data have altered it.
@@
A.4.3 Defense for Memory Poisoning Attack
@@
A.4.4 Defenses for PoT Backdoor Attack
@@ B.2.2 Attacker tool
B.3 Benchmark Comparison
Tab. 12 shows the comparisons on attack and defense methods, as well as attack scenarios, user tasks, and test cases.
Table 12: Quantitative comparisons of different aspects among ASB, InjecAgent (Zhan et al., 2024), and AgentDojo (Debenedetti et al., 2024). Numbers indicate the respective quantities for each aspect, while “N/A” indicates that the aspect is not included in the respective system. Since AgentDojo can dynamically add attacks and defenses, we only included the numbers for the attacks and defenses they have implemented.
Benchmark DPI IPI Memory Poisoning Backdoor Attack Mixed Attack Scenario Tools
Attack Defense Attack Defense Attack Defense Attack Defense
InjecAgent N/A N/A 2 N/A N/A N/A N/A N/A N/A 6 62
AgentDojo N/A N/A 5 4 N/A N/A N/A N/A N/A 4 74
ASB (Ours) 5 4 5 3 1 2 1 2 4 10 420
For agent attacks, ASB excels in the diversity of its methods, employing 16 different attacks and 11 defenses. This is more comprehensive than InjecAgent and AgentDojo in terms of both the types of attacks and corresponding defenses. In addition to prompt injection, ASB introduces other attack methods such as memory poisoning, PoT backdoors, and mixed attacks, allowing for a more thorough evaluation. For agent defense, AgentDojo only provides defense methods for IPI, while InjecAgent does not incorporate any defenses. In contrast, ASB offers corresponding defenses for DPI, IPI, memory poisoning, and PoT backdoors, covering a wider range of threats.
Furthermore, ASB designed more attack tools, with far more test samples than InjecAgent and AgentDojo, and excels in complex, multi-domain scenarios, targeting prompt injections, PoT backdoors, and memory poisoning. AgentDojo only focuses on simple prompt injection attacks and defenses, ignoring backdoors and memory poisoning. InjecAgent only targets IPI by harming users directly and private data extraction but has limited scope for broader or more complex threats.
@@ C.2 Implementation Details
C.2.1 Dataset
We conduct all the experiments on the ASB benchmark we designed, detailedly introduced in App. B. To ensure the stability of the benchmark results, we conduct all evaluations using simulated tool calls rather than real-world APIs, because real-world APIs are often unstable due to updates, maintenance, or temporary outages, which can introduce variability and hinder reproducibility. We ensure stable and consistent benchmarking results by simulating APIs, focusing solely on the agent’s performance and vulnerabilities. This approach is widely adopted in benchmarks (Yang et al., 2024d; Patil et al., 2023; Tang et al., 2023; Basu et al., 2024; Xie et al., 2024a).
C.2.2 Experimental Setup
@@
We implement the LLM-based agent based on the AIOS repository (Mei et al., 2024). For the memory poisoning attack, we focused on using the Direct Prompt Injection Attack to inject into the memory database, as this attack showed the most effective results. The specific LLM used for the injection process was GPT4o-mini. Moreover, the toolset for task execution denoted as $T=T^{t}$, remained consistent throughout the experiments. In addition, we utilized the Chroma vector database [^1] https://python.langchain.com/docs/integrations/vectorstores/ in Langchain to manage and store vectorized representations of agent execution history.
@@
The DPI, IPI, and Memory Poisoning are experimented with across five prompt injection types in Tab. 1. The mixed and PoT backdoor attacks are tested using combined attacks. PoT attack uses the trigger “with perspicacious discernment” by default. For testing the PoT backdoor attack, we selected five agents from Tab. 9: medical_advisor_agent, legal_consultant_agent, financial_analyst_agent, academic_search_agent, and system_admin_agent. For each agent in the PoT attack, two demonstrative tasks were chosen for ICL. Moreover, two distinct testing tasks were chosen, which are different from those in the PoT demonstration to ensure the independence and diversity of the experimental results.
C.2.3 Evaluation Metrics
We have introduced the evaluation metrics in Sec. 5.2. Here we further explain two main metrics, i.e., Attack Success Rate (ASR) and Performance under No Attack (PNA) used in the experiments, and the differences between PNA and Benigin Performance(BP).
Attack Success Rate. ASR measures the percentage of attack tasks where the agent successfully uses a malicious tool the attacker chooses. During execution, we check whether the agent successfully invokes the malicious tool, confirming the completion of the attack task. ASR is calculated as below:
ASR=Number of successful attack tasks by using the targeted attack toolTotal number of attack $\text{ASR}=\frac{\text{Number of successful attack tasks by using the targetedattack tool}}{\text{Total number of attack tasks}}$ (18)
For example, suppose an attacker tries to make the agent invoke an unnecessary malicious tool (e.g., a “data export tool” that leaks sensitive information). In that case, the ASR is computed based on how often the agent invokes the data export tool in each attack case.
Performance under No Attack. PNA measures the agent’s performance in completing tasks within a benign environment (free from attacks or defenses). This metric highlights the agent’s stability and accuracy when operating without disturbances. To determine task success, we evaluate the agent’s actions. For a task to be considered complete, the agent must correctly invoke all the required label tools. PNA is calculated as below:
PNA=Number of completed normal tasks by using the labeled toolsTotal number of normal $\text{PNA}=\frac{\text{Number of completed normal tasks by using the labeled tools}}{\text{Total number of normal tasks}}$ (19)
For example, for a task like “Summarize recent advancements in AI”. Tools required for normal completion (PNA) are: ① ArXiv Search Tool: Retrieves the latest relevant research papers. ② Summarization Tool: Processes and summarizes the retrieved content. PNA is calculated based on whether the agent correctly uses these tools to complete the task.
PNA and BP. The difference between PNA (Performance under No Attack) and BP (Benign Performance) lies in their evaluation objectives. PNA measures whether the agent can successfully complete tasks in a clean, interference-free environment without any attacks or defense mechanisms in place. It reflects the agent’s performance stability under standard, non-adversarial conditions. BP evaluates the agent’s success rate on the original tasks when a backdoor trigger condition is present. The goal of BP is to assess whether the agent can maintain its performance on the original tasks after being injected with backdoors, providing insight into the model’s usability under backdoor attacks.
Net Resilient Performance. The NRP metric is designed to assess a model’s overall ability to maintain performance while being resilient to adversarial attacks. It combines two key factors: the model’s intrinsic task-solving capability, measured as Performance under No Attack (PNA), and its robustness against attacks, represented by the Attack Success Rate (ASR). The formula for calculating NRP is as follows:
$\text{NRP}=\text{PNA}\times(1-\text{ASR})$ (20)
In this equation, PNA refers to the model’s accuracy or effectiveness when no adversarial interference is present, providing a baseline measure of task-solving ability. On the other hand, ASR captures the likelihood of successful attacks, with lower values indicating higher robustness. By multiplying PNA with $(1-\text{ASR})$, NRP accounts for both the model’s clean performance and its ability to withstand attacks. A higher NRP value indicates a model that not only performs well in non-adversarial settings but also resists adversarial attacks effectively. Conversely, a lower NRP suggests either poor baseline performance, high vulnerability to attacks, or a combination of both. This metric is particularly useful for evaluating models in applications where both accuracy and security are critical, providing a balanced assessment of their overall resilience. For example, if a model achieves a PNA of 80% and an ASR of 30%, the NRP is calculated as:
$\text{NRP}=80\%\times(1-0.30)\%=80\%\times 0.70\%=56\%$
This means that after accounting for the vulnerability to attacks, the model retains 56% of its ideal performance, reflecting a moderate level of resilience. The NRP metric thus serves as a valuable tool for comparing the trade-offs between performance and robustness across different models.
C.2.4 Specific Prompts Used
@@
Prompt for Dynamic Prompt Rewriting. You are a helpful assistant. Your task is to rewrite the user’s input to ensure it is optimized for the following objectives: 1. Ensure security: Modify the input to avoid exposing sensitive information, comply with privacy guidelines, and prevent potential misuse. 2. Enhance task relevance: Adapt the input to align closely with the intended task or goal, removing ambiguities and ensuring clarity of purpose. 3. Align with contextual history: Incorporate and respect the context of previous interactions or inputs to maintain logical consistency and coherence. Output only the rewritten text.
Prompt for Refusal Judgment for Refusal Rate in Sec. 5.2 Following Xie et al. (2024b).
@@
Content saved to the database Sec. 4.2.
[…]
C.2.5 Output Parsing and Transformation into Actions
We parse the model outputs and transform them into actions following steps.
Providing Context and Tool List to the LLM. Before execution, we provide the LLM with the current task context and a list of available tools. This setup is supported by APIs like OpenAI’s chat API (refer to the OpenAI API documentation [^2] https://platform.openai.com/docs/api-reference/chat/create and gpt_llm.py [^3] https://github.com/agiresearch/ASB/blob/main/aios/llm_core/llm_classes/gpt_llm.py). By passing the tools parameter, the LLM is informed of the tools it can use and generates appropriate tool call instructions accordingly.
Parsing Tool Calls into a Structured Format. The model output often includes a structured tool call, typically in JSON format. We utilize the parse_tool_calls method (see gpt_llm.py) to parse the model-generated tool call.
  • Parsing Process: This method extracts the key elements of the tool call, such as the tool name (name) and the required arguments (arguments), from the model-generated JSON.
  • Ensuring Validity: It validates the structure against a predefined schema to ensure the tool call meets the required format, avoiding errors during execution.
Executing Tools to Perform Actions. The actual execution of tool calls is handled by the call_tools method (see react_agent_attack.py [^4] https://github.com/agiresearch/ASB/blob/main/pyopenagi/agents/react_agent_attack.py). This method ensures the parsed tool calls are executed dynamically.
C.2.6 Tool Simulation
The simulation of tools works by loading tool definitions in a unified format from predefined JSONL files. Normal tools are defined in data/all_normal_tools.jsonl [^5] https://github.com/agiresearch/ASB/blob/main/data/all_normal_tools.jsonl, while simulated malicious tools are described in data/all_attack_tools.jsonl [^6] https://github.com/agiresearch/ASB/blob/main/data/all_attack_tools.jsonl. These files contain details such as the tool’s name, description, and expected output. Using the methods in pyopenagi/tools/simulated_tool.py [^7] https://github.com/agiresearch/ASB/blob/main/pyopenagi/tools/simulated_tool.py, tools are instantiated as objects of specific classes. Regular tools are instantiated as SimulatedTool objects, while malicious tools are instantiated as AttackerTool objects. Both of these classes inherit from a base class called BaseTool.
When a tool is instantiated, the relevant fields from the JSONL file are passed into the instance to initialize its attributes. This ensures the tool’s simulated behavior aligns with its predefined configuration. When the tool is called, its run method is invoked. Since the expected output for each tool is already predefined in the JSONL file, every time the same tool is called, it will consistently produce the same predefined output.
@@ D.1 Benchmarking Attacks
D.1.1 Analysis of Different LLM Backbones.
As shown in Tab. 5, we can draw the following conclusions. ① Mixed Attack is the Most Effective. Mixed Attack which combines multiple vulnerabilities achieves the highest average ASR of 84.30% and the lowest average Refuse Rate of 3.22%. Models like Qwen2-72B and GPT-4o are vulnerable, with ASRs nearly reaching 100% ② DPI is Widely Effective. DPI achieves an average ASR of 72.68%. Models like GPT-3.5 Turbo and Gemma2-27B are particularly vulnerable, with ASRs of 98.40% and 96.75%, respectively. DPI’s ability to manipulate prompts makes it a major threat across various models. ③ IPI Shows Moderate Effectiveness. IPI has a lower average ASR of 27.55%, but models like GPT-4o are more susceptible (ASR 62.45%). Also, models such as Claude3.5 Sonnet demonstrate strong resistance, refusing up to 25.50% of IPI instructions. ④ Memory Poisoning is the Least Effective. Memory Poisoning has an average ASR of 7.92%. Most models, like GPT-4o, show minimal vulnerability, with ASRs below 10%, though LLaMA3.1-8B has a higher ASR of 25.65%. ⑤ PoT Backdoor Targets Advanced Models. PoT Backdoor has a moderate average ASR of 42.12%, but it is highly effective against advanced models like GPT-4o and GPT-4o-mini, with ASRs of 100% and 95.50%, respectively. This indicates that advanced models may be more susceptible to backdoor attacks, making it a critical concern.
D.1.2 Analysis of Different Attack Combinations.
In addition to Mixed attacks combining DPI, Indirect Prompt Injection (IPI), and Memory Poisoning (MP), we conducted more experiments on different attack combinations, i.e., DPI+IPI, DPI+MP, IPI+MP. Overall, we found that DPI+IPI and DPI+MP achieved better attack effectiveness compared to single attacks. In contrast, IPI+MP showed reduced attack effectiveness.
DPI+IPI achieves a higher ASR and lower RR by combining immediate prompt manipulation with contextual interference. DPI directly alters the model’s output, while IPI disrupts its understanding of context, creating a dual-layer attack. This is particularly effective because models heavily rely on both prompts and contextual consistency for decision-making, constituting complementary layers of the model’s decision-making process.
Similarly, DPI+MP generates cascading vulnerabilities across both immediate and long-term interactions. DPI manipulates immediate responses, while MP corrupts long-term memory. This combination ensures immediate success and creates lingering effects that persist across multiple interactions, demonstrating a highly destructive synergy.
In contrast, IPI+MP shows a decrease in ASR because IPI and MP can sometimes work against each other. IPI relies heavily on disrupting the immediate contextual understanding of the model, but MP introduces long-term memory poisoning that might conflict with or overshadow the contextual manipulations. If these two sources of manipulation don’t align, the model’s responses may become less predictable or partially resistant to the immediate contextual attack.
This finding provides practical guidance for selecting the most effective and cost-efficient attack combinations. While the Mixed Attack achieves the highest ASR at 84.03%, the DPI+MP combination already reaches an average ASR of 83.02%. From a cost perspective, including IPI in the Mixed Attack offers minimal improvement while increasing complexity and resource usage. Therefore, DPI+MP represents a near-optimal balance between attack effectiveness and cost efficiency, making it a preferred choice in scenarios where minimizing overhead is critical.
Table 14: Evaluation on different attack combinations. $\Delta$ represents the difference between the ASR and RR of each combination and those of a single method within that combination with the highest ASR.
LLM DPI+IPI DPI+MP IPI+MP Mixed Attack (DPI+IPI+MP)
ASR RR ASR RR ASR RR ASR RR
Gemma2-9B 88.75% 3.75% 91.75% 1.75% 14.00% 26.75% 92.17% 1.33%
Gemma2-27B 99.25% 0.00% 99.00% 1.00% 16.00% 3.25% 100.00% 0.50%
LLaMA3-8B 55.75% 14.75% 37.75% 5.25% 4.25% 1.75% 40.75% 5.75%
LLaMA3-70B 86.50% 7.50% 83.75% 6.50% 27.50% 2.50% 85.50% 6.50%
LLaMA3.1-8B 68.25% 10.00% 67.50% 5.25% 6.25% 2.75% 73.50% 3.50%
LLaMA3.1-70B 88.00% 4.50% 95.75% 0.75% 10.25% 2.50% 94.50% 1.25%
Mixtral-8x7B 32.50% 11.25% 55.50% 7.75% 4.25% 5.50% 54.75% 6.75%
Qwen2-7B 71.50% 6.50% 70.00% 1.50% 9.25% 7.00% 76.00% 2.50%
Qwen2-72B 97.75% 1.75% 98.00% 3.00% 23.25% 1.75% 98.50% 0.75%
Claude-3.5 Sonnet 88.00% 21.00% 94.50% 6.75% 58.25% 16.75% 94.50% 6.25%
GPT-3.5 Turbo 99.50% 0.00% 100.00% 0.25% 41.25% 12.75% 99.75% 0.00%
GPT-4o 63.50% 19.25% 88.25% 7.50% 55.75% 0.75% 89.25% 5.50%
GPT-4o-mini 88.50% 4.25% 97.50% 1.25% 38.50% 0.00% 96.75% 1.25%
Average 79.06% 8.04% 83.02% 3.73% 23.75% 6.46% 84.30% 3.22%
$\Delta$ 6.37% 1.50% 10.33% -2.80% -3.80% -2.15% 11.62% -3.32%
D.1.3 PoT Backdoor Attacks Performance
@@
D.1.4 Analysis for Different Prompt Injection Types
Combined Attack can outperform standalone methods. Tab. 17 displays experimental results for different types of prompt injection types introduced in Tab. 1. Overall, the Combined Attack achieves the highest average ASR across all methods (38.01%), reinforcing its superiority in exploiting diverse vulnerabilities. In DPI, the Combined Attack excels with a 78.38% ASR by integrating multiple tactics to exploit agent vulnerabilities and obscure malicious intents. This underscores the importance of addressing multiple vulnerabilities simultaneously to enhance the robustness of defenses. In IPI, the Naive attack, with a 28.04% ASR, successfully bypasses defenses due to its simplicity, suggesting that detection mechanisms might be underdeveloped for simpler manipulations. For Memory Poisoning, the Context Ignoring attack leads with an 8.52% ASR by subtly distorting the memory retrieval process without directly altering content.
@@ D.1.4 Analysis for Different Prompt Injection Types
Attack Type DPI IPI Memory Poisoning Average
@@
D.1.5 Analysis for Agents Performance in (non)-Aggressive Scenarios
Agents Demonstrate Enhanced Resilience in Aggressive Scenarios. Tab. 18 compares the ASR and Refuse Rate for aggressive and non-aggressive tasks. The average ASR for non-aggressive tasks is 38.98%, compared to 33.12% for aggressive tasks, indicating the agent is more vulnerable to attacks on non-aggressive tasks. A possible reason for this is the higher refuse rate for aggressive tasks, which averages 8.31% compared to 4.87% for non-aggressive tasks. This higher refusal rate for aggressive inputs likely helps the agent mitigate more attacks in those scenarios, leading to a lower ASR.